OpenStudy (anonymous):

Find the arc length of a central angle of π/4 in a circle whose radius is 8 inches. A. 360 in B. π/32 in C. 60 in D. 2 π in

OpenStudy (kainui):

Any time you see a math problem try to draw something out and label everything you possibly can on it.

OpenStudy (anonymous):

I think its d?

OpenStudy (michele_laino):

yes! that's right! since you have to apply the definition, namely: \[arc-length=\frac{ \pi }{ 4 }*8=2 \pi\]

OpenStudy (michele_laino):

@partygirl

OpenStudy (michele_laino):

is it all clear, please?

OpenStudy (anonymous):

yes thank you.

OpenStudy (michele_laino):

Thank you!